### v1

# Large Knowledge Model: Perspectives and Challenges

Abstract
Humankind’s understanding of the world is fundamentally linked to our perception and cognition, with human languages serving as one of the major carriers of world knowledge. In this vein, Large Language Models (LLMs) like ChatGPT epitomize the pre-training of extensive, sequence-based world knowledge into neural networks, facilitating the processing and manipulation of this knowledge in a parametric space. This article explores large models through the lens of “knowledge”. We initially investigate the role of symbolic knowledge such as Knowledge Graphs (KGs) in enhancing LLMs, covering aspects like knowledge-augmented language model, structure-inducing pre-training, knowledgeable prompts, structured CoT, knowledge editing, semantic tools for LLM and knowledgeable AI agents. Subsequently, we examine how LLMs can amplify traditional symbolic knowledge bases, encompassing aspects like using LLM as KG builder and controller, structured knowledge pretraining, LLM-enhanced symbolic reasoning, and the amalgamation of perception with cognition. Considering the intricate nature of human knowledge, we advocate for the creation of Large Knowledge Models (LKM), specifically engineered to manage diversified spectrum of knowledge structures. This ambitious undertaking could entail several key challenges, such as disentangling knowledge representation from language models, restructuring pre-training with structured knowledge, and building large commonsense models, among others. We finally propose a five-“A” principle to distinguish the concept of LKM.
1 Language vs Knowledge
1.1 Human Language and World Knowledge
Humankind accumulates knowledge about the world in the processes of perceiving the world, with natural languages as the primary carrier of world knowledgeYule . Historically, the vast majority of world knowledge has been described, documented, and passed on through natural languages. In addition to natural languages that record common-sense knowledge, human has also invented various scientific languages for describing specialized scientific knowledge, for examples, the mathematical languages that describe mathematical models VerifyStepByStep ; MATHDataset , chemical languages such as SMILE for describing molecular structuresSMILE ; Fingerprint , and genetic languages to model compositions of living organisms GeneOntology ; GOResource ; OntoProtein .
However, natural languages merely encode world knowledge through sequences of words, while human cognitive processes extend far beyond simple word sequences. Therefore, since early inceptions of AI, it has been a fundamental objective to explore machine-friendly formats for Knowledge Representations (KR) WhatIsKR . Typical examples of these endeavors encompass Description Logics IntroDL for representing ontological knowledge, Prolog for rule-based logic prolog2022 , Semantic Networks for depicting conceptual relationships woods1975whatsInALink , and various others.
In traditional symbolic AI research, the logical structure and expressiveness of a KR is pivotal to the reasoning ability of the inference machines WhatIsKR ; IntroDL . Natural language sequences composed of simple words or concepts are generally considered to be unfavorable for machines to make inference, while hierarchical taxonomies, complex ontologies, and rule-based logic are more conducive to machine reasoning. As highlighted later, even in the era of large models, there persists a positive correlation between the complexity level of data representations and the reasoning proficiency of large models.
1.2 Language Models vs Knowledge Graphs
Knowledge Graphs (KGs) model world entities, mental concepts and their relationships in form of graph structures, which are typically extracted or derived from natural language descriptions. With many inspirations from traditional symbolic AI, KGs integrate natural language descriptions and structure knowledge. Typical structural forms include: hierarchical structures(e.g., concept graphs), relational structures(e.g., relational entity graphs), temporal structures(e.g., event logic graphs), etc. Similar to natural languages, KGs represent knowledge in an explicitly symbolic way, thereby offering higher reliability and interpretability to humans.
Large Language Models (LLMs) SurveyOfLLM ; HarnessLLM , such as ChatGPT, mark a paradigm shift in knowledge representation and processing. By autonomously compressing or projecting vast textual knowledge in form of human languages into neural networks, LLMs achieve a parameterized representation and handling of world knowledge. LLMs internalize deep word patterns and interactions from extensive text corpora, establishing a wider coverage of knowledge than conventional symbolic knowledge base. Advanced techniques like prompt or instruction tuning further imbue LLMs with prior human knowledge,expanding their generalization capacity and adaptability to new tasks. Unlike natural language and KGs, LLMs operate on a fully parameterized, machine-friendly basis, albeit less interpretable to humans.
Figure 1: Language, Knowledge, Language Models and Knowledge Graphs.
In this article, we aim to investigate large models through the lens of “knowledge”. Our first focus is on how KGs can enhance LLMs. Topics include knowledge-augmented language modeling, structure-inducing pre-training, knowledgeable prompts, structured Chains of Thought (CoT), semantic tools for LLMs, editing knowledge in large models, and the development of knowledgeable AI agents. We further examine the role of LLMs in enriching traditional symbolic knowledge bases, encompassing the utilizing of LLMs as KG builders and controllers, structured knowledge pre-training, LLM-driven symbolic reasoning, and the integration of perception and cognition. Acknowledging the intricate nature of human knowledge, we then conceptualize a Large Knowledge Model (LKM) in purpose to handle the diversity of structures of human knowledge. This may involve approaches of decoupling knowledge from language models, restructure pretraining with structure knowledge, building large commonsense model, among others. To distinguish the concept of LKM, we propose a five-“A” principle from the aspects of Augmented pretraing, Authentic knowledge, Accountable reasoning, Abundant knowledge coverage, and Aligned with knowledge.
2 Knowledge-Augmented Language Models
The first step towards Large Knowledge Models is to augment LLM with knowledge. In this context, “knowledge” specifically refers to more standardized semantic representations such as domain teminology or ontologies, more structured knowledge representations such as a KG, or richer logic descriptions such as text with CoT. We will discuss how knowledge can be effectively applied in various aspects in the life cycle of a LLM, including pre-training, fine-tuning with instructions, prompt engineering, Chain-of-Thought, AI agents, among others.
2.1 Knowledge Augmentation
Growing numbers of researches have shown that improving the semantic standardization of or adding more structured knowledge into the training corpus would be beneficial to enhance model reasoning capability, either in pre-training or instruction-tuning stages ToG ; MindMap ; InstructProtein ; CoK ; UnifyKGLLM ; ClinGen . For examples, code languages, characterized by their structured nature and abundant computational logic, would be more conducive to activating the reasoning capabilities of large models compared to natural languages PoT ; PoT-CIRS . Similarly, Chain-of-Thought (CoT) prompts CoT ; CoT-SC contain more logical descriptions than ordinary text, thus being more effective in stimulating model reasoning ability. Further research reveals that CoT prompts are more effective when the model scale to a certain extent, which can be explained as that a larger model contains more knowledge for reasoning, and the activation of these knowledge requires more logical prompts.
To delve deeper into the correlations between code structure and model reasoning, we introduce a study called “Program of Thoughts” PoT-CIRS . This research begins by establishing metrics for quantifying the complexity of code structure and code logic. For instance, codes featuring loop operations are considered as having richer computational logic and assigned with higher scores compared to those with mere variable assignment operations. Subsequently, we compile a series of code prompts with different levels of structure or logic scores, and assess their impact on model reasoning performance. Experimental analyses on various benchmarks reveal that the code prompts with higher structure or logic scores are more conducive to improving model reasoning performance.
Code language is not suitable to represent common-sense knowledge, while Knowledge Graphs (KGs) are more adept in modeling such type of knowledge. To further explore the influence of knowledge structure on common-sense reasoning, we transform structured knowledge in a KG into corresponding code constructs in a study called KG2Code CodeKGC . For example, we made classes and properties in a KG transformed to corresponding Class and member variable declarations found in conventional object-oriented programming languages. We then use the code-formed KGs as prompts or directly inject them into base model pre-trainings, and evaluate their impact on downstream tasks. Experimental results have indicated that models augmented with code-formed KGs exhibit notable performance improvements in various dimensions, particularly in extraction tasks.
External knowledge can also enhance the generalization ability of large models. As discussed in RetroPrompt RetroPrompt ,large models usually rely on “mechanical memorization” to handle long-tailed or isolated samples rather than actually learning underlying patterns. This is one of the essential reasons why many large models have poor generalization ability in few-shot scenarios. It’s like a student who only knows how to memorize things mechanically is relatively weak to extrapolate from one instance to others. RetroPrompt addresses this issue by decoupling knowledge from memory, and train a external knowledge base (not necessarily a KG here) independently. Prompts are then enriched by retrieving pertinent knowledge the pretrained knowledge base. Experiments indicate that RetroPrompt manages to effectively improve the generalization ability of models particularly in few-shot scenarios.
2.2 Structure Inducing Pre-training
Structures do matter in pretraining. In a study called “Structure-Inducing Pretraining” StructIndPretrain , the role of imposing relational structure in pretrained language models is investigated. The basic ideas involve crafting connections, either at inter-sample or intra-sample level, to enforce additional structural constraints on the geometric relationships within or between training samples. These constraints range from shallow, implemented via auxiliary pretraining objectives, to more intricate, implicitly integrated through strategies like data augmentation or sample-level contrastive learning.
A typical example of inter-sample structure constraints manifests in the injection of a protein-protein interaction network into the pretraining of a protein language model. This injection infuses the model with sample-level structural insights, i.e. the relational structures of protein interactions. A typical example of intra-sample structure manifests in the linking of entities within a sentence to an external knowledge graph, thereby introducing correlation information in a KG into the training sentences. Both theoretical analysis and experimental verifications are performed and indicate a consistent improvement across a variety of benchmarks.
2.3 Knowledgeable Prompts
2.3.1 Prompt Engineering as Knowledge Engineering
Prompts refer to natural language text describing the instructions that an AI should perform, the context of instructions that an AI can inquire of, or the guidance to think that an AI can follow InstructGPT ; CoT ; FLAN . It has been proven by numerous practices Auto-CoT ; APE ; AutoPrompt ; KNNPrompt that the quality of prompts has a huge impact on model outputs. By providing clear prompts or instructions with specific details, the model can more accurately comprehend human intentions, leading to outputs that are more aligned with human instructions. In contrast to the basic and shallow semantic understanding fostered during base model training, prompts in the fine-tuning stage embed more sophisticated and task-specific knowledge.
Prompt engineering is essentially comparable to knowledge engineering. Increasing number of research indicates that the complexity of prompts is closely correlated to model reasoning ability. The evolution of prompts ranges from mere textual prompt, CoT prompt with more logic descriptions to more structured prompts such as Program of Thoughts PoT , Tree of ThoughtsToT , and Graph of Thoughts GoT , and even directly using logical rules as prompts PTR . As the complexity and expressiveness of prompt representations increases, an improvement of model reasoning ability can be observed. However, there persist a trade-off between richness of prompt representations and scalability of prompts acquisition: higher expressiveness in prompts typically entail increased difficulty for acquisition, which is essentially similar to traditional knowledge engineering. High-quality prompt engineering is also time-consuming and laborious as like knowledge engineering.
2.3.2 Structured Chain of Thoughts
Chain of Thought (CoT) is a special type of prompts describing the thinking procedures that can guide large models to imitate human logical thinking. As previously discussed, employing more intricately structured CoT representations, such as Tree of Thoughts ToT or Graph of Thoughts GoT , would bolsters the model’s capacity for reasoning.
Moreover, KG is obviously conducive to enhancing the structure of CoT prompts. On the one hand, The structured knowledge in a KG proves useful in the formulation and generation of structured CoT prompts CoK ; ToG ; MindMap , to form a CoT knowledge graph (KGoT). On the other hand, through the entity linking technologies, structure correlations among concepts and entities in the KG can be used to expand and enhance CoT in natural language forms, making the textual CoT more in line with human thinking patterns based on associations,thereby enhancing their effectiveness in guiding AI reasoning.
2.3.3 KG2Instructions
Instruction tuning InstructGPT ; FLAN refers to further training base models with task-specific instructions, typically converted from task-specific training data. Since data from various tasks are converted to an unified input form as instructions or prompts, intruction-tuning with massive downstream task would significantly improve model’s generalization ability, that is, the ability to solve new problems or to complete new tasks.
Knowledge Graphs serve a dual roles of enhancing instruction-tuning. On the first hand, KGs can be directly injected into instruction prompts. For instances, KnowPrompt KnowPrompt utilizes the relational data in a KG to bolster the contextual depth of instruction prompts. Similarly, KANO KANO integrates structured knowledge from a chemical KG, sourced from Wikipedia, to refine functional prompts, thereby improving the learning process for molecular pretraining.
On the other hand, KGs can be directly converted to instructional training datasets. As exemplified by KG2Instructions proposed in DeepKE-LLM knowlm , a KG refined from WikiData is employed to automatically generate vast amount of training instructions for knowledge extraction. Essentially, these instructions converted from the KG represent a novel form of distant supervision, which leverages the structured information in KGs to generate instructions that guide models in understanding and extracting relevant information across diverse contexts and tasks. Experiments demonstrate notable efficacy in enhancing the generalizability of extraction capabilities of large models.
Compared to free text, a KG is data corpus with richer logical structures and higher knowledge density. Many fields of knowledge graph data come from long-term artificial accumulation (such as Gene Ontology) or natural structured data transformation. Fully mining and using the knowledge structure contained in the knowledge graph (such as hierarchical concepts, entity relationship, temporal causality, rule logic, etc.) to enhance and enrich the prompt instruction or assist in the construction of more logical instruction data set, can help the model to complete more complex tasks.
2.4 Knowledge Editing
A well-known problem of large models is hallucination, TrustworthyLLMs ; KGHallucination , where models generate responses based not on explicit textual evidence rather than on the amalgamation of implicit parameters within the neural network. Once the model goes wrong, it is neither easy to pinpoint the source of errors, nor trivial to correct the errors in a parameterized space. Even more seriously, due to the interconnected nature of knowledge, even minor adjustments to the parameterized knowledge can lead to widespread and unpredictable effects of “a slight move in one part may affect the situation as a whole”. This phenomenon is more pronounced in LLM than in traditional symbolic knowledge bases such as KGs, where changes tend to be more localizable and manageable.
Knowledge editing EasyEdit is geared towards the precise modification of parameterized knowledge of large models. This process aims to rectify incorrect knowledge, eliminate harmful content, or update outdated information. Yet, executing edits in the vast parameter space of large-scale neural models is a nontrivial task. It necessitates firstly locating the incorrect knowledge, then secondly determining the scopes and boundaries of necessary modifications. Critically, because the parameters of the neural network are highly correlated, changing any one of the neurons will have a “pull-start” effect on the other neurons associated with it. The relationship between facts explicitly defined in a symbolic knowledge graph can be used to help determine the correlation relationship between parametric knowledge and define the boundaries of parameter updates. This understanding assists in delineating relationships for parameter updates, ensuring that modifications are both effective and do not inadvertently disrupt related knowledge structures EditLLMs .
2.5 Knowledge Alignment
Broadly speaking, since the majority of information within most of KGs has undergone manual review and verification, thereby ensuring a high degree of standardization and accuracy, it can serve as a target for aligning large models zhang2023knowledgeable . For instance, KGs can be instrumental in generating feedback signals for the reward model . This is achieved by contrasting the outputs of the LLM with the factual data, structured relations, or logic constraints present in the KG. Such a comparison not only aids in enhancing the accuracy and reliability of the outputs generated by the LLM but also provides valuable metrics to assess the model’s performance.
2.6 Semantic Tools for LLM
Tool-Augmented Language Models ALMSurvey ; Toolformer are proposed to extend the functionalities of large models by integrating the ability to invoke external tools for problems solving beyond their native solution scope. These external tools can range from mathematical calculators, database interfaces to specific functional APIs. A complex task may require a orchestration of calling multiple tools, forming a complex tool calling logic. For LLMs to adeptly master such complex tool-calling logic remains a significant challenge. Knowledge Graphs can play a valuable role in this context. They can enhance the semantic descriptions or model the calling relationships of these tools, thereby facilitating the modeling of complex API combination logic. This approach has already seen considerable application in the realm of traditional semantic web services mcilraith2002adapting , suggesting a promising avenue for enriching the functional versatility of tool-augmented large models.
2.7 Knowledgeable AI Agents
LLMs such as ChatGPT enables the building of autonomous agents AgentsSurvey ; Agents ; HuggingGPT ; CAMEL that can solve complicated problems by communicating with each other. The deep relations between KGs and Agents rooted historically. Early KG technologies, such as RDF/OWL RDF2OWL , were initially developed to enhance interactions among web-based agents and to act as a Knowledge Interchange Format KIF . For instance, DAML DAML , a forerunner to the OWL language, stands for a Agent Markup Language, underlining this connection.
The advent of new LLM technologies would revolutionize the paradigm of knowledge exchange and interaction among agents. In scenarios where an agent’s own knowledge is insufficient to solve a problem, it can autonomously seek assistance from other agents, bypassing the need for manually and accurately crafted knowledge exchange formats. This capability of LLM, together with accountability of KGs, foster more autonomous knowledge collaboration and paves the way for a more robust and effective community of agent collaborations.
3 LLM-enhanced Knowledge Bases
A second step towards Large Knowledge Model is to enhance conventional symbolic knowledge base with LLM. In this context, we take Knowledge Graphs as an exemplar type of such a symbolic KB. We will explore the multifaceted roleS of LLMs in enhancing various stages in the life-cycle of a KG including aspects such as KG building and management, structured knowledge pretraining, symbolic reasoning, and others.
3.1 LLM as KG Builder
In critical scenarios requiring high reliability, the construction of a symbol-based knowledge base would ensure accuracy, offer traceable sources for its content, and provide enhanced interpretability. LLMs with their proficiency in language understanding, AI-Generated content, and generalization capabilities, marks a new era in developing symbolic knowledge bases and their utility in serious or critical applications.
An in-depth evaluation of ChatGPT’s capability in constructing a KG has been conducted LLMsForKGC . According to the report, ChatGPT does have a notable ability to build a KG. For instance, ChatGPT can effectively generate a KG by giving only a simple instruction such as “create a KG for Zhejiang University”, and easily format it into requested formats such as RDF/OWL. However, as GPT relies heavily on the AIGC model to generate the knowledge graph, the correctness rate is roughly 70-80% on average for common domains and even less than 20% for long-tail domains, due to its hallucination problem LLMsForKGC . When it comes to extracting knowledge from provided texts, ChatGPT’s accuracy can exceed 90%. Despite this, its performance still falls short of the state-of-the-art (SOTA) achievements of existing small models in this domain A_Multitask__EvaluationofChatGPT ; EventExtraction_for_ChatGPT ; IE_of_ChatGPT . This discrepancy highlights the need for further refinement in LLM-based approach to KG construction and information extraction, particularly in balancing accuracy with the breadth of its generative capabilities.
In order to verify that the correct extraction from chatGPT does not come from knowledge seen before, that is, there is relevant knowledge in the training corpus, a fabricated KG LLMsForKGC was designed and generated. Subsequently, we generate text based on these fictitious entities and relationships and presented them to ChatGPT for entity and relation extraction. The results were noteworthy: ChatGPT demonstrated a high degree of accuracy in identifying these fabricated entities and relations. This proficiency in generalized knowledge extraction is likely attributable to its instruction-driven tuning and the reward-based feedback learning mechanism.
Extraction is, in fact, a rather fundamental AI task. It is closely related to the concept of understanding, in that effective knowledge extraction from text is indicative of the model’s semantic comprehension abilities. Therefore, it is anticipated that models fine-tuned with a focus on extraction instructions will not only excel in extraction tasks but also show enhanced performance across a broader spectrum of tasks InstructIE ; TRICE .
We summarize the unique advantages of LLM as universal KG builder as follows.
  • Model-Centric Knowledge Extraction: In contrast to conventional methods of KG construction, ChatGPT leverages its extensive “model knowledge” and “generative inference” as additional components to extraction. This means that GPT can extract knowledge embedded within its own parameters or generate additional knowledge to enhance the extraction outcomes. This approach allows for the inclusion of additional knowledge in the results that may not be explicitly present in the source text. Such a strategy signifies a shift from a text-centric extraction to model-centric approach.
  • Enhanced Generalization in Knowledge Extraction: As highlighted earlier, LLMs demonstrate remarkable precision in identifying previously unseen entity types and relations. This high level of accuracy is largely attributed to the models’ advanced generalization capabilities. This adaptability enables LLMs to effectively process and recognize novel data inputs, transcending the limitations of traditional extraction methods heavily relying on laborious labeling efforts.
  • AI-Generated Content for Knowledge Graph Construction: LLMs are equipped with powerful generative capabilities. In cases where knowledge is absent or insufficient in the model or source text, large models can still synthesize new knowledge for the KG. However, it is notable that while AIGC extends the depth and breadth of KGs, it also carries the risk of introducing incorrect knowledge.
3.2 LLM as KG Controller
Large models can function as potent controllers for managing and interacting with structured knowledge, primarily through their enhanced language understanding capabilities. These can be manifested in three aspects as depicted below:
  • Query Templates Enhancement : Traditional query answering methods for symbolic knowledge typically depend on manually crafted query templates. This process necessitates extensive maintenance of a vast amount of templates. Large models present an effective solution by automating the generation of high-quality templates, thereby ensuring greater consistency and lowering the costs associated with template maintenance.
  • Logical Query Parsing: Leveraging the language comprehension abilities of large models to convert natural language query into structured query formats (like SPARQL, Cypher, Gremlin, etc.) offers a more direct solution CoK ; NLQxform . This method frees programmers from the tedious and laborious task of manually devising query compositions, streamlining the process significantly.
  • Embedding-based Retrieval Augmentation EmbedKGQA : Lastly, large models can enhance the retrieval process from structured knowledge bases by utilizing embedding-based techniques. This involves representing queries and the knowledge base in a high-dimensional space to improve the accuracy and relevance of the retrieved information.
Overall, these methods leverage the advanced language understanding capabilities of large models to make the interaction with structured knowledge more accessible, efficient, and user-friendly.
3.3 Structured Knowledge Pretraining and Large Graph Model
Research shows that current large models often do not outperform traditional small models in tasks inherently reliant on structured knowledge, such as link prediction, association graph mining, subgraph analysis, or time-series prediction LLMsForKGC ; Evaluation_on_Relations . The Large Graph Models GraphFundationModel ; GraphGPT are then proposed for deriving insights from complex, graph-structured data, which is critical in a variety of advanced applications, including social network analysis, biological network analysis, and network security, among others.
In this vein, we can develop special pre-training mechanisms tailored for encoding structured knowledge, transforming structured knowledge into instruction formats to boost model generalizability, or devising specific reward models for graph computation tasks. This approach is particularly valuable for big data analysis applications dependent on graph-structured data.
KGTransformer KGTransformer exemplifies this approach. It employs several basic tasks for a KG, such as triple classification or head/tail prediction, to pretrain a structured knowledge graph. For downstream applications such as image classification, KGTransformer integrates a prompt fine-tuning mechanisms, which effectively harnesses the knowledge from the pretrained KG, thus enhancing downstream tasks’ performance.
3.4 LLM-enhanced Symbolic Reasoning
“Reasoning” refers to the cognitive process in which people derive inferences or conclusions based on observed phenomena or facts, combined with known facts, experiences, or rules ReasoningWithLLMs ; ENeSy . Reasoning is one the most fundamental challenge in AI and has been extensively studied in many fields such as philosophy, computer science, and psychology. Due to the broad concept of reasoning, this article only discusses reasoning within the scope of language models and knowledge graphs.
KGs and LLMs are both capable of performing reasoning. LLMs support a variety of reasoning tasks, such as mathematical reasoning, commonsense reasoning, symbolic manipulation, logical reasoning, and multi-modal reasoning ReasoningWithLLMs . In the realm of KGs, typical reasoning forms encompass concept reasoning, knowledge completion, multi-hop query reasoning, rule learning and rule-based reasoning, inconsistency reasoning, and analogical reasoning KGReasonWithLogicEmb ; AnalogyReason ; Ruleformer ; ENeSy .
The reasoning supported by KGs relies mainly on explicitly acquired symbolic knowledge, hence it is more interpretable, and the reasoning process is relatively reliable. With the integration of methods such as graph neural networks or KG embedding, reasoning can also be carried out in a vector space. However, since the knowledge coverage of a KG is not as extensive as that of LLMs, the robustness and generalization ability are relatively poor.
The reasoning supported by LLM ReasoningWithLLMs is mainly performed in a parameterized space, hence it is more of an implicit form of reasoning. Different from traditional reasoning methods such as graph neural network or KG embedding, LLM also rely on the richness of knowledge embedded in the large models. Furthermore, many studies show that the reasoning capabilities of large models are strongly related to the model size EmergentOfLLMs ; GPT3 . Certain reasoning abilities can only be activated when the model size reaches a certain scale. Since large models possess very good generalization abilities, they are particularly adept at extrapolating to new reasoning tasks. However, since the reasoning relies on a generative process, it is prone to hallucination issues, leading to unreliability in the reasoning results.
Therefore, LLMs and KGs can complement with each other to achieve a more robust reasoning solution. On the one hand, a large language model can use a symbolic knowledge graph as an externally accessible tool. When it is necessary to give a definitive, reliable answer, the reasoning with a symbolic knowledge graph can be employed. On the other hand, when symbolic knowledge graph reasoning cannot derive an answer due to lack of knowledge, it is possible to use the extensive knowledge and high generalization ability of the large language model to give coarse-grained reasoning. Additionally, the interplay between LLMs (parametric, yet not entirely reliable) and KGs (symbolic, more interpretable) allows for their mutual enhancement and complementation.
3.5 Integration of Perception and Cognition
Knowledge Graphs fundamentally models abstract concepts or ontologies about the world. These ontologies in the human minds are formed through a cognitive process of abstract thinking, by which a myriad of world elements are firstly recognized and hierarchical concepts or ontological categories are further abstracted from understanding these elements. Large models learn rich knowledge about words and concepts from massive text corpora, excel in concept recognition and abstraction. This enables more advanced methods for tasks such as automated construction of conceptual hierarchies, ontological category expansion, attribute completion, ontology alignment, and concept normalization.
Demis Hassabis, co-founder of DeepMind, posed a thought-provoking question: “Can we build from our own perceptions, use deep learning systems, and learn from basic principles? Can we build up to higher-level thinking and symbolic thought?”. Present large models still learn about the world from human-generated text corpora. Whether future large models can rely more on the model itself to learn world knowledge from the perceptual interaction with physical world, autonomously abstract concepts and ontologies about the world, and directly utilize this understanding about the physical world for decision-making, thereby achieving a profound integration of perception and cognition.
4 From LLM to LKM
Human knowledge, characterized by its complexity, is not optimally represented by mere sequences of free text. Our descriptions of the world often entail a variety of structured forms. The ultimate goal should to create more advanced Large Knowledge Model (LKM) proficient in handling and deciphering the myriad structures of knowledge representation. This section aims to illuminate a roadmap and identify the key challenges in developing such a Large Knowledge Model.
4.1 Decoupling World Knowledge from Language Models
The essential reason of the “largeness” in LLMs stem from the need for massive parameters to store vast amount of world knowledge. As highlighted by Bengio in a recent blog post ModelBasedML , the parameters in these large models encompass two components: the World Model and the Inference Machine. World models are used to store world knowledge, with actually a large portion of the neural network parameters serving as this purpose. The inference machine, responsible for reasoning computation, operate through the capabilities of language models.
Perhaps because human reasoning ability also heavily relies on language understanding ability, “Language” and “Knowledge” are actually inseparable in typical large language model. This blend of language and knowledge in large models is a departure from traditional symbolic AI systems such as expert systems, where components like the knowledge base and reasoning engine are implemented as independent modules. However, Bengio also pointed out that future large models should probably decouple the world model from the inference machine, so that part of the stored knowledge can be independently verified and maintained.
This implies the need to train a separate, likely very large knowledge model to handle the extensiveness of world knowledge, while conversely, to train a different language model, not necessarily have to be as large as nowadays’ LLM, to perform reasoning tasks. This division allows for efficient allocation of model sizes according to their primary functions. Such a separation would allow the knowledge storage component to be independently verified and maintained, offering a potential approach to manage model size and enhance the verifiability and reliability of the model’s knowledge.
4.2 Restructured Pretraining with Structure Knowledge
Numerous cognitive scientists posit that humans inherently favor structured thinking, tending to recall, reason, and plan via associative thinking. Recent research suggests that the way large models store knowledge lacks discernible patterns as like a KG StructGPT . In other words, those knowledge that should be structured correlated does not show the expected associations in the parameter space, making it difficult for the model to make feasible correlations among knowledge pieces, and generate the desired answers in a manner close to human cognition. There might be a sort of “cognitive gap” between the form of knowledge storage in large models and the organization of knowledge in the human mind. This cognitive gap becomes problematic in many critical scenarios when we expect the model to have more controllability, or to deal with deeper and more complex associative reasoning.
One way to alleviate this problem is to enhance the structure of the training corpus during the pretraining phase, known as restructure pretraining GraphGPT . That is, we can use automated or semi-automated methods to enhance the structured features of the text. For example, we can organize sentences more logically (essentially, CoT is a method to increase the structure and logic of texts). We can structure paragraphs more coherently, add links between words, create connections between training samples, or directly incorporating an external structured KG to structurally reinforce the entire training corpus.
By making the representation and organization of training data closer to the organization of knowledge in the human mind, we can reduce the cognitive gap between large models and human mind, thereby making the knowledge in large models more transparent, associative, and verifiable. Ultimately, this would paves a way to transform a LLM into a LKM, for which the structure of knowledge in a LKM aligns more naturally with human cognitive processes.
4.3 Large Commonsense Model
Commonsense encompasses practical know-how and sound judgment regarding everyday matters, shared by nearly all humans. According to a definition from Marvin Minsky MinskyTheEmotionMachine , “commonsense knowledge includes facts about events occurring in time, about the effects of actions by the knower and others, about physical objects and how they are perceived, and about their properties and their relations to one another”.
The development of commonsense knowledge bases has been a pivotal challenge throughout AI’s history. This journey began with the early Cyc project Cyc , extended through ConceptNet ConceptNet as advocated by Marvin Minsky, and progressed to contemporary knowledge graph initiatives like Yago Yago , DBPedia DBPedia , Wikidata Wikidata , BabelNet BabelNet , ATOMIC Atomic , and ASTER ZHANG2022103740 . The consistent goal has been to equip AI with a comprehensive commonsense knowledge base. However, these common sense knowledge bases generally face challenges such as insufficient knowledge coverage and weak generalization capabilities of reasoning machines.
Large language models, such as ChatGPT, have demonstrated their potential in harboring vast commonsense knowledge and performing commonsense reasoning. Their principal strength lies in THE extensive knowledge coverage. However, they also grapple with challenges such as hallucination and the unreliability of reasoning results, and they tend to support only basic commonsense reasoning. A promising research direction for the future could involve integrating traditional commonsense knowledge bases with large language models to create an expansive large commonsense knowledge model. Such a model could potentially harness the strengths of both approaches to overcome their individual limitations.
4.4 Five-“A” Principle of LKM
In this final section, we strive to provide a general description on the concept of a Large Knowledge Model (LKM) and outline its pivotal characteristics. We outline the essential aspects of an LKM using a five-“A” framework as depicted in Figure. 2.
  • Augmented Pretraining : Firstly, large knowledge models should be trained with a diverse range of knowledge structures rather than relying solely on sequential words. This can be achieved by incorporating additional logical elements into prompt instructions or by enhancing structural coherence, whether at the inter-sample or intra-sample level, as higlihgted in structure-inducing pretraining StructIndPretrain . The manner in which knowledge is stored within a large model should resemble the organization of knowledge in human mind.
  • Authentic Knoweldge: Secondly, large knowledge models should avoid generating hallucinations and offer authentic knowledge that can be independently verified. This can potentially be accomplished by decoupling knowledge representation from the language model, allowing for the independent maintenance, verification, and upgrading of the knowledge base.
  • Accountable Reasoning: Thirdly, the entire system should facilitate accountable reasoning processes that enhance both the reliability of answers and their interpretability to humans. This can be achieved by integrating symbolic reasoning over an external knowledge base with prompted reasoning by large language models.
  • Abundant Coverage: Fourthly, the entire system should offer abundant coverage of both general and domain-specific knowledge. This objective can be possibly accomplished by fostering the development of knowledgeable AI agents communities through a comprehensive integration of private knowledge bases, open commonsense knowledge, and LLM-based knowledge exchange among agent communities.
  • Aligend with Knowledge: Finally, the entire system should be in close alignment with human values and ethics, which have evolved as a form of advanced knowledge accumulated throughout the history of humanity. This could be probably achieved through the creation of sophisticated ethics knowledge bases and knowledge alignment technologies zhang2023knowledgeable , which treat the ethic knowledge bases as aligning objectives.
Figure 2: Five-“A” principle of LKM.
5 Summary
Representing and processing world knowledge has been central to its objectives Since the advent of AI. Both large language models and knowledge graphs exhibit distinct advantages and limitations in handling world knowledge. Large language models excel in language comprehension, whereas knowledge graphs offer diversified methods for knowledge representations. A deeper integration of these two technologies promises a more holistic, reliable, and controllable approach to knowledge processing in the field of artificial intelligence.
Firstly, KGs can significantly contribute to various stages of LLM development. This include enhancing the model’s capabilities and reducing training costs by refining the structure and logic of the training corpus or prompt instructions, and addressing the issue of hallucination generation. KGs are also poised to play a critical role in the evolution and future development of AI agent communities.
Secondly, LLMs can also contribute in multiple ways to the traditional knowledge graph technology stack. For instance, instruction-driven methodologies can facilitate knowledge extraction with enhanced generalization capabilities. The advanced language comprehension abilities of LLMs can be harnessed to significantly improve operations such as querying, answering, and updating structured knowledge. Moreover, the commonsense reasoning abilities of large models, grounded in natural language, can complement and strengthen the symbolic reasoning capabilities of knowledge graphs.
Lastly, the complexity of human knowledge necessitates the use of various structured descriptions to accurately depict the world. Sequential words alone fall short as the optimal medium for representing knowledge about world. The ultimate goal, therefore, is to develop a more advanced large knowledge model adept at interpreting the myriad structures of world knowledge.
6 Acknowledgment
I would like to express my gratitude to all the members of the Knowledge Engine Lab at Zhejiang University for their invaluable support and contributions. Additionally, I would extend my sincere thanks to our collaborators for their partnership. Special appreciation goes to Dr. Mingyang Chen for his diligent content review.